OpenStudy (anonymous):

write 3x+2y=14 in slope intercept form

OpenStudy (anonymous):

y=-3/2x + 7

OpenStudy (anonymous):

slope intercept form: y = mx+b so 3x+2y =14 - subtract 3x to both sides 2y=-3x+14-- divide 2 to both sides y = (-3x+14)/2-- simplify y = -3x/2+7